OpenStudy (aworldofcolor):

I don't understand how to solve this question and I was wondering if someone will help. I'll give medals (: evaluate the expression for the given value of the variable 8a + 6;q =2

OpenStudy (aworldofcolor):

I forgot to add the options which are 10 14 20 22

OpenStudy (dayakar):

8a+6 =8*2+6 (a=2) =16+6

OpenStudy (dayakar):

add it